OpenStudy (jinx.exe):

What is the approximate distance around the half circle with a diameter of 64 meters?

OpenStudy (anonymous):

The distance around a full circle is the circumference: \[c = \pi d\] where d is the diameter. So the distance around half the circle is just half that, or: \[32\pi \approx 3.14159 \cdot 32 \approx 100.53\]

OpenStudy (jinx.exe):

so whats the answer @adotm ;-;?

OpenStudy (anonymous):

The last equation is the half circumference, both in exact form (32 pi) and approximate decimal (100.53).

OpenStudy (jinx.exe):

A. 100.48 m B. 200.96 m C. 401.92 m D. 1607.68 m these are the only answers @adotm

OpenStudy (anonymous):

It's A, which is: \[3.14 \cdot 32\] I just used a more exact number for pi.

OpenStudy (jinx.exe):

okay